1. Use the method of implicit differentiation to determine the derivatives of the following:


a. xsin y + ysin x =1

b. tan(x-y) =y/1+x^2

C. √x+y = x^4 + y^4

d.y + xcos y = x^2 y


2. Find the number "c" that satisfy the Mean Value Theorem (M.V.T) on the given intervals:


a. f(x)= e^-x ; [ 0,2 ]

b. f(x)=x/x+2 ; [ 1, π ]


3.Determine the equation of the tangent and normal at the given points:


a. y + xcos y = x^2 y ; [ 1, π/2 ]

b. h(x) = 2/√x^2 + 1 ; at x=1



1
Expert's answer
2022-01-10T15:41:37-0500

1.

a.


"x\sin y + y\sin x =1"

Diffirentiate both sides with respect to "x"


"\sin y+(x\cos y)y'+(\sin x)y'+y\cos x=0"

"y'=-\dfrac{\sin y+y\cos x}{\sin x+x\cos y}"

b.


"\tan(x-y) =\dfrac{y}{1+x^2}"

"(1+x^2)\tan(x-y)=y"

Diffirentiate both sides with respect to "x"


"2x\tan(x-y)+\dfrac{(1+x^2)(1-y')}{\cos ^2(x-y)}=y'"

"y'=\dfrac{2x\tan(x-y)+1+x^2}{1+x^2+\cos ^2(x-y)}"

c.


"\sqrt{x+y} = x^4 + y^4"

Diffirentiate both sides with respect to "x"


"\dfrac{(1+y')}{2\sqrt{x+y}}=4x^3+4y^3y'"

"y'=\dfrac{8x^3\sqrt{x+y}-1}{1-8y^3\sqrt{x+y}}"

d.


"y + x\cos y = x^2 y"

Diffirentiate both sides with respect to "x"


"y'+\cos y-(x\sin y)y'=2xy+x^2 y'"

"y'=\dfrac{2xy-\cos y}{1-x\sin y-x^2}"

2.

Find the number "c" that satisfy the Mean Value Theorem (M.V.T) on the given intervals:

a. "f(x)=e^{-x}, [0, 2]"


"f'(x)=-e^{-x}"

"f'(c)=-e^{-c}, c\in(0, 2)"

"f(0)=e^{-0}=1"

"f(2)=e^{-2}"

Use the Mean Value Theorem


"f'(c)=\dfrac{f(2)-f(0)}{2-0}"

"-e^{-c}=\dfrac{e^{-2}-1}{2-0}"

"e^c=\dfrac{2}{1-e^{-2}}"

"c=\ln(\dfrac{2}{1-e^{-2}})"

b. "f(x)=\dfrac{x}{x+2}, [1, \pi]"

"f'(x)=\dfrac{x+2-x}{(x+2)^2}=\dfrac{2}{(x+2)^2}"

"f'(c)=\dfrac{2}{(c+2)^2}, c\in(1, \pi)"

"f(1)=\dfrac{1}{1+2}=\dfrac{1}{3}"

"f(\pi)=\dfrac{\pi}{\pi+2}"

Use the Mean Value Theorem


"f'(c)=\dfrac{f(\pi)-f(1)}{\pi-1}"

"\dfrac{2}{(c+2)^2}=\dfrac{\dfrac{\pi}{\pi+2}-\dfrac{1}{3}}{\pi-1}, c\in(1, \pi)"

"\dfrac{2}{(c+2)^2}=\dfrac{3\pi-\pi-2}{3(\pi-1)(\pi+2)}"

"\dfrac{1}{(c+2)^2}=\dfrac{1}{3(\pi+2)}"


Since "c\in (1, \pi)," then


"c=\sqrt{3(\pi+2)}-2"

3.Determine the equation of the tangent and normal at the given points:

a.


"y'+\cos y-(x\sin y)y'=2xy+x^2y'"

"y'=\dfrac{\cos y-2xy}{x\sin y+x^2-1}"

Point "(1, \pi/2)"


"slope_1=m_1=\dfrac{\cos (\pi/2)-2(1)(\pi/2)}{1\sin (\pi/2)+(1)^2-1}=-\pi"

The equation of the tangent line in point-slope form


"y-\pi/2=-\pi(x-1)"

The equation of the tangent line in slope-intercept form


"y=-\pi x+3\pi/2"

"slope_2=m_2=-\dfrac{1}{m_1}=\dfrac{1}{\pi}"

The equation of the normal line in point-slope form


"y-\pi/2=\dfrac{1}{\pi}(x-1)"

The equation of the normal line in slope-intercept form


"y=\dfrac{1}{\pi} x+\pi/2-\dfrac{1}{\pi}"

b.


"h'(x)=-\dfrac{2x}{(x^2+1)^{3/2}}"

"x=1"


"slope_1=m_1=h'(1)=-\dfrac{2(1)}{(1^2+1)^{3/2}}=-\dfrac{\sqrt{2}}{2}"

"h(1)=\dfrac{2}{\sqrt{1^2+1}}=\sqrt{2}"

The equation of the tangent line in point-slope form


"y-\sqrt{2}=-\dfrac{\sqrt{2}}{2}(x-1)"

The equation of the tangent line in slope-intercept form


"y=-\dfrac{\sqrt{2}}{2} x+\dfrac{3\sqrt{2}}{2}"

"slope_2=m_2=-\dfrac{1}{m_1}=\sqrt{2}"

The equation of the normal line in point-slope form


"y-\sqrt{2}=\sqrt{2}(x-1)"

The equation of the normal line in slope-intercept form


"y=\sqrt{2}x"